OpenStudy (anonymous):

Prove Sn: 2 is a factor of n^2 - n + 2

OpenStudy (anonymous):

What I thought would work is do \[S _{k}:2x =k ^{2}-k+2\] \[S _{k+1}:2x =(k+1) ^{2}-(k+1)+2\] Then \[S _{k} = S _{k+1}\]

OpenStudy (anonymous):

To me, I would go on this way if n is odd, then n = 2k +1 for some k in R then n^2 = (2k+1)^2 = 4k^2+4k +1 therefore \(n^2 -n +2 = 4k^2+4k+1 -2k-1+2=4k^2+2k+2=2(k^2+k+1)\) hence, n^2 -n +2 divided by2 If n is even, then n = 2m n^2 = 4m^2 \(n^2 -n+2 = 4m^2 -2m +2 = 2(2m^2-m +1)\) that shows n^2 -n+2 divided by 2 also. If n =0, then 2|2 All cases lead to the proof done.

ganeshie8 (ganeshie8):

If you want to work it by induction, you need to start by showing that the given statement is true for base case @darielgames

ganeshie8 (ganeshie8):

\[S _{k+1}: (k+1) ^{2}-(k+1)+2 = (k^2 - k +2) + 2k\]

ganeshie8 (ganeshie8):

From induction hypotheisis \(\large k^2 -k+2 = 2x\) so \[S _{k+1}: (k+1) ^{2}-(k+1)+2 = (k^2 - k +2) + 2k = 2x + 2 = 2y\] ends the proof

OpenStudy (anonymous):

I understand what you did right here (k+1)2−(k+1)+2=(k2−k+2) but why do you add 2k? and where exactly does 2 and 2y come in?

ganeshie8 (ganeshie8):

expand \(\large (k+1)^{2}-(k+1)+2 \)

ganeshie8 (ganeshie8):

\(\large (k+1)^{2}-(k+1)+2 = k^2 + 2k + 1 - k - 1 + 2\) yes ?

OpenStudy (anonymous):

Alright then the 2 = 2y, you are using y for \[S_{k}\]?

OpenStudy (anonymous):

yes thats right Sorry just making it linear is a bit congusing

OpenStudy (anonymous):

confusing

ganeshie8 (ganeshie8):

from induction hypothesis (Sk) we know that stuff inside parenthesis is a multiple of 2 right ?

ganeshie8 (ganeshie8):

\[\large \begin{align} (k+1)^{2}-(k+1)+2 &= k^2 + 2k + 1 - k - 1 + 2 \\~\\&= (k^2 - k +2) + 2k \\~\\ &= (2x) + 2k\end{align}\]

OpenStudy (anonymous):

Sorry I don't understand >.> What you did was simplification or equality in this one ? http://puu.sh/bvxIZ/161657d03d.png

OpenStudy (anonymous):

Omg.... I get it!

OpenStudy (anonymous):

k^2 - k +2 is the base so it's the same thing as saying 2x

OpenStudy (anonymous):

@ganeshie8 Thanks!

ganeshie8 (ganeshie8):

yw :) btw, don't forget to prove the statement for base case n=0